OpenStudy (anonymous):

A container has 50 electronic components, of which 10 are defective. If 6 components are drawn at random from the container, what is the probability that at least 4 are good?

OpenStudy (anonymous):

P(at least 4) = P(4) + P(5) + P(6) P(4) = (6C4) (40/50) (39/49) (38/48) (37/47) (10/46) (9/45) P(5) = (6C5) (40/50) (39/49) (38/48) (37/47) (36/46) (10/45) P(6) = (40/50) (39/49) (38/48) (37/47) (36/46) (35/45) add them all up and it's around 0.9164

OpenStudy (anonymous):

i meant 0.9144